The only force acting on a 5kg object has components Fx=20N and Ft=30N. Find the magnitude and directions of the acceleration of the object.


1
Expert's answer
2021-01-31T03:41:52-0500

"a_x=F_x\/m=20\/5=4(m\/s^2)"


"a_y=F_y\/m=30\/5=6(m\/s^2)"


"a=\sqrt{4^2+6^2}=7.21(m/s^2)" . Answer


"\tan\alpha=6/4=1.5\to \alpha=56.3°" from the positive direction of the x-axis. Answer